• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ole Hansen

    Log entry time 17:07:23 on April23,2003

    Entry number 98307

    keyword=Shift Summary

    8:00 Start shift with optics data taking in progress (run 1233) at p=1.1932 (+4%),
    raster on.
    8:35 End run 1233. Start run 1234.
    9:07 End run 1234. Start run 1235 (last run of this set).
    9:35 End run 1235. Moved target to multi-foil C. Start run 1236.
    9:50 Current increased to 5uA (was 1uA). Start run 1237.
    10:06 End 1237, start 1238.
    10:20 End 1238 (2M). Asking for raster off. Start 1239.
    10:32 End 1239. Beam goes to Hall C for BCM calibrations.
    10:34 Changing momentum to 1.1817 (+3%) and septum current to 102.047A.
    10:39 Start 1240 at new kinematics, raster off, multi-foil target.
    10:40 Beam goes away for controlled access, as requested by out RC.
    Hall C doing current calibrations meanwhile.
    12:15 Controlled access will take significantly more time because leak on
    reference cell is hard to fix.
    13:40 Controlled access finished. Beam returns. Start run 1241.
    13:45-57 Beam trips frequently (unacceptable)
    14:10 Noticed Q2R tripped. End run 1241 (end probably garbage).
    14:25 Q2R reset, cycled, and restored. Start run 1242.
    14:30 We are informed that the IOC controlling Q1R/L apparently has crashed.
    Current readback frozen, no heatbeat. Field readback still alive and ok,
    so probably no data lost. Trying to reset IOC from upstairs fails.